OpenStudy (anonymous):

A box contains 1 green, 2 yellow, and 3 red marbles. Two marbles are drawn at random without replacement. What are the odds of drawing two different colors?

OpenStudy (anonymous):

this is not that easy i don't think plus we have to remember that you are asked for "odds' not "probability"

OpenStudy (anonymous):

probably easiest to figure out the probability of getting two of the SAME color

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

could be two yellow that would be \[\frac{2}{6}\times\frac{1}{5}=\frac{1}{15}\]

OpenStudy (anonymous):

So that's yellow... But there's only 1 green, so I can't have two of those...

OpenStudy (anonymous):

of could be two red

OpenStudy (anonymous):

that would be \[\frac{3}{\binom{6}{2}}=\frac{3}{15}\]

OpenStudy (anonymous):

add and get \[\frac{4}{15}\] as the probability of getting two of the same color so \[\frac{11}{15}\] of getting different colors but that is not the answer

OpenStudy (anonymous):

because you are asked for "odds" not 'probability' 11 out of 15 times get different colors 4 out of 15 get the same odds are 11:4

OpenStudy (anonymous):

Oh, I get it! Thank you so much :)

OpenStudy (anonymous):

yw